# Mixed-state geometric phases of coherent and squeezed spin states

### v2

@@ IV.1.2 The interferometric geometric phase
To study the IGP of the SSS, we consider the unitary evolution
$\rho(\Theta(t))=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta(t))}=S(\Theta(t))\rho(\Theta(0))S^{\dagger}(\Theta(t)),$ (65)
where $\rho(\Theta(0))=\rho(0)=\frac{1}{Z}\mathrm{e}^{-\beta H}$. To ensure that $S$ satisfies the parallel-transport condition (18), it is required that
00 0 $=\operatorname{Tr}\left[\rho(0)S^{\dagger}\dot{S}\right]$
$=\sum_{m=-j}^{j}\left\langle jm\right|\frac{\mathrm{e}^{-\beta\omega_{0}J_{z}}}{Z}\frac{2J_{-}J_{+}+2J_{z}}{8}\left|jm\right\rangle\dot{\Theta}$
$=\sum_{m=-j}^{j}\frac{\mathrm{e}^{-m\beta\omega_{0}}}{Z}\frac{j^{2}+j-m^{2}}{4}\dot{\Theta}.$ (66)
However, this condition cannot be satisfied at finite temperatures in general, so the IGP in its original form is not guaranteed. The parallel-transport condition (18) and its strengthened form (19) are both satisfied if $\dot{\Theta}=0$. However, this implies that $\Theta$ is a constant, such that $S(\Theta(t))=S(\Theta(0))=1$ for any $t$, resulting in a trivial evolution. Therefore, the IGP of the one-axis SSS also becomes trivial due to the parallel-transport condition.
[…]
Nevertheless, one can obtain the non-adiabatic IGP based on Eq. (22) accumulated during an arbitrary evolution beyond the parallel-transport condition (18) or the strengthened version (19). To fairly compare with the results of the Uhlmann phase of the SSS, in the following we choose $j=1$ for the non-adiabatic IGP as well. The total phase of the $j=1$ SSS is given by
$\theta_{\text{T}}(t)=\arg\operatorname{Tr}[\rho(0)S(\Theta(t))]$
$=$ $\arg\left[\left(\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta(t)}{2}}\right)\left(\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}\right)+\mathrm{e}^{-\frac{\mathrm{i}\Theta(t)}{2}}\right],$ (67)
and the non-geometric phase is
@@
$\theta_{\text{NG}}(t)$ $=-\mathrm{i}\int_{0}^{t}\operatorname{Tr}[\rho(0)S^{\dagger}(t^{\prime})\dot{S}(t^{\prime})]\mathrm{d}t^{\prime}$
@@
$=-\frac{2+\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}}{1+\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}}\frac{\Theta(t)}{4}.$ (68)
The non-adiabatic IGP is then obtained by $\theta^{\prime}_{\text{G}}(t)=\theta_{\text{T}}(t)-\theta_{\text{NG}}(t)$. Since the evolution in this case is not necessarily periodic, $\theta_{\text{G}}$ also depends on the final (instantaneous) time $t$.
In the top panel of Figure 6, the non-adiabatic IGP $\theta_{\text{G}}^{\prime}$ is plotted as a function of temperature $T$ and the squeezing parameter $\Theta_{f}\equiv\Theta(t_{\text{final}})$. Different from the Uhlmann phase of the $j=1$ SSS, there is no temperature-induced geometric phase transition here. We show the behavior of $\theta^{\prime}_{\text{G}}$ as a function of $T$ for $\Theta_{f}=5.4$ in the bottom panel, in which $\theta^{\prime}_{\text{G}}$ continuously decreases with $T$. At finite temperatures, we find that $\theta^{\prime}_{\text{G}}$ also continuously varies with $\Theta_{f}$ without any discontinuities.
However, in the zero-temperature limit, $\theta_{\text{T}}=\arg[1+\mathrm{e}^{-\frac{\mathrm{i}\Theta_{f}}{2}}]$ while $\theta_{\text{NG}}$ is always a continuous function of $\beta$ and $\Theta_{f}$. Therefore, as $\Theta_{f}$ crosses $2\pi$, the midpoint of the parameter range $[0,4\pi]$, the value of $\theta^{\prime}_{\text{G}}$ experiences a $\pi$-jump (from $\frac{\pi}{2}$ to $-\frac{\pi}{2}$ or vice versa), indicating a geometric phase transition. We show this behavior of $\theta^{\prime}_{\text{G}}$ in the middle-left panel of Figure 6. For comparison, we also plot $\theta^{\prime}_{\text{G}}$ vs $\Theta_{f}$ at $T=1.0\omega_{0}$ in the middle-right panel to contrast the smoothing of the curve at finite temperature. For the $j=1$ SSS, therefore, finite-temperature jumps of the two types of mixed-state geometric phases are absent.
@@ IV.2 Two-Axis Squeezing
Although realizing a nontrivial parallel transport process for the IGP via one-axis squeezing for the SSS is difficult, here we will show that it is possible in the two-axis squeezing scenario. In the following, we will investigate two-axis squeezing for the Uhlmann phase and IGP.
@@ IV.2.1 Uhlmann Phase
Figure 7: Uhlmann phase $\theta_{U}$ of two-axis squeezing of the $j=1$ SSS as a function of temperature $T$ on semi-log scale. The smooth curve indicates the absence of a transition.
@@
Figure 8: (Top panel) The IGP $\theta_{\text{G}}$ of the $j=1$ two-axis SSS as a function of $T$ and $\theta_{f}$, where $\theta_{f}\in[0,\frac{3\pi}{4}]$. $\theta_{\text{G}}$ is continuous with respect to both $\theta_{f}$ and $T$ (apart from a jump at $T=0$). (Bottom panels) $\theta_{\text{G}}$ vs $T$ at $\theta_{f}=\pi/2$, where $\theta_{\text{G}}$ exhibits no jumps (left), and $\theta_{\text{G}}$ vs $\theta_{f}$ at $T=0$, where there is a $\pi$-jump at $\theta_{f}\approx 1.33\,\text{rad}$ (right).
@@ IV.2.2 The IGP
where all diagonal elements are zero. Therefore, the parallel transport condition  (19) is equivalent to
@@
Figure 9: Contour plot of the non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ for the two-axis SSS with $j=1$ as a function of temperature $T$ and the squeezing parameter $\theta_{f}$, where $\theta_{f}=\phi_{f}$ corresponding to $\theta_{f}\in[0,\frac{3\pi}{4}]$.
Similar to the previous examples, the non-adiabatic IGP of two-axis squeezing can be expressed as
$\theta^{\prime}_{\text{G}}(t)=\arg\operatorname{Tr}[\rho(0)K(t)]+\mathrm{i}\int_{0}^{t}\operatorname{Tr}\left[\rho(0)K^{\dagger}(t^{\prime})\dot{K}(t^{\prime})\right]\mathrm{d}t^{\prime}.$ (92)
The evolution is characterized by a simple curve $\kappa(t):\theta(t)=\phi(t)$ with $\theta_{f}\in[0,\frac{3\pi}{4}]$ selected in the parameter space. This curve starts at $\theta=0$ when $t=0$ and ends at any point $(\theta_{f},\phi_{f})$ along the path, where $\theta_{f}\equiv\theta(t_{\text{final}})$ and $\phi_{f}\equiv\phi(t_{\text{final}})$. To contrast with the IGP, we present the contour plot of $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\theta_{f}$ for the $j=1$ two-axis SSS. The overall configuration of this plot demonstrates remarkable similarity to Fig. 8. This is expected given that the second term in Eq. (92) represents a continuous function $g(\theta_{f},T)$ dependent on $\theta_{f}$ and temperature $T$. As shown in Figure 9, the non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ remains continuous as a function of $T$, independent of the specific value of $\theta_{f}$.
Nevertheless, there is a $\pi$-jump of the non-adiabatic IGP at $\theta_{f}=1.33$ rad when $T=0$, similar to the IGP shown in Fig. 8. The evolution of the IGP, however, is generated by the path with constant $\phi$ and respects the parallel-transport condition. The discontinuity stems solely from the first term of Eq. (92), a feature that remains invariant irrespective of the parallel-transport condition. Consequently, the zero-temperature jump survives in the non-adiabatic IGP as well. These findings support feasibility for experimentally probing geometric phase transitions of mixed states. Even when evolution strictly adhering to the parallel-transport condition is experimentally challenging, the framework of non-adiabatic IGP may still characterize possible geometric properties of the system.
@@ V Implications
Since the CSS and SSS have been realized in condensed matter or atomic, molecular, and optical systems Estève et al. (2008); Gross et al. (2010); Huang et al. (2021); Byrnes (2024); Hosten et al. (2016), experimental generation and measurement of the mixed-state geometric phases may become feasible in the future. However, imposing the parallel-transport condition for the IGP or Uhlmann phase requires precise controls of the systems and the ancilla representing the environment, which can be challenging in laboratories. Nevertheless, the non-adiabatic IGP relaxes the constraint and becomes a prominent candidates for demonstrating interesting geometric effects of CSS and SSS at finite temperatures. We caution that the non-adiabatic IGP is buried in the total phase, and a systematic removal of the non-geometric contribution according to Eq. (22) is required to obtain the genuine geometric contribution.
On the other hand, the rapid development of quantum computers opens possibilities of simulating complex quantum systems. By using an ancilla qubit acting as the environment and coupling it to a system qubit, Ref. Viyuela et al. (2018) shows that the Uhlmann phase of a spin-$1/2$ system can be simulated on a quantum computer by using two engineered Hamiltonians to respectively evolve the system and ancilla simultaneously. Since a spin-$j$ object may be decomposed as a collection of $2j$ spin-$1/2$ objects, one may use $2j$ qubits as a system of a spin-$j$ system coupled to an ancilla system of $2j$ qubits to realize the CSS or SSS coupled to the environment. It then follows a series of technical challenges to construct suitable engineered Hamiltonians for the system and ancilla in order to generate the corresponding IGP or Uhlmann phase when the corresponding parallel-transport condition is imposed during the evolution. After the targeted geometric phase is generated on a quantum computer, one may follow Ref. Viyuela et al. (2018) to couple additional measurement qubits to the composite system and perform state tomography to extract the phase accumulated during the constrained evolution.
@@ VI Conclusion
We have derived the Uhlmann phase and the IGP along with its non-adiabatic generalization for the CSS and SSS. Selected examples are fully solved to demonstrate interesting properties of those mixed-state geometric phases. The Uhlmann phase is well-defined for both CSS and SSS due to its rigorous mathematical structure, but there may or many not be topological phase transitions manifested by discrete jumps of the Uhlmann phase as temperature varies. In contrast, the parallel-transport condition for the IGP only renders a trivial solution for the one-axis SSS although the IGP of the CSS exhibits discrete jumps. The non-adiabatic IGP forgoes the constraint and exhibits discrete jumps in the geometric part of the accumulated phase for the CSS. However, no finite-temperature jump is found in the non-adiabatic IGP for both one-axis and two-axis squeezing with $j=1$. Our formalisms provide the foundation for future theoretical and experimental investigations of mixed-state topological properties of quantum spin systems.
@@ ACKNOWLEDGEMENTS
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X.Y.H. was supported by the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656.
@@ Appendix A Correspondence between Uhlmann phase and Berry phase of CSS as $T\rightarrow 0$
where $J_{z}|j,-j\rangle=-j|j,-j\rangle$ has been applied. This result precisely reduces to the Berry phase of the ground state $|j,-j\rangle$ Chaturvedi et al. (1987). More specifically, when $j=\frac{3}{2}$ and $\theta=\frac{\pi}{2}$, $|\zeta|=1$ and thus $\theta_{U}=\mathrm{i}\frac{3}{4}{\oint}_{\gamma}(-2\mathrm{i}\tan^{2}\frac{\theta}{2})\mathrm{d}\phi=3\pi\equiv\pi\mod 2\pi$, which confirms the result shown in Fig. 1 as $T\rightarrow 0$.

### v3

@@ IV.1.2 The interferometric geometric phase
As shown in the discussion around Eq. (II.4), the operator $\tilde{S}$ automatically satisfies the strengthened parallel-transport condition of the IGP given by Eq. (20). Accordingly, the IGP accumulated during this evolution is given by
@@
$\theta_{\text{G}}(t)=\arg\operatorname{Tr}[\rho(0)\tilde{S}(t)]=\arg\left(\sum_{m}\lambda_{m}\nu_{m}\mathrm{e}^{\mathrm{i}\phi_{m}}\right),$ (64)
where $\lambda_{m}$ is the $m$-th eigenvalue of the density matrix, $\nu_{m}=\langle m(0)|m(t)\rangle$ is the overlap between the initial and final eigenstates, and
$\phi_{m}$ $=\mathrm{i}\int^{t}_{0}\mathrm{d}t^{\prime}\langle jm,\Theta(t^{\prime})|\frac{\mathrm{d}}{\mathrm{d}t^{\prime}}|jm,\Theta(t^{\prime})\rangle$
$=\mathrm{i}\int^{t}_{0}\mathrm{d}t^{\prime}\langle jm|S^{\dagger}(t^{\prime})\frac{\mathrm{d}S(t^{\prime})}{\mathrm{d}t^{\prime}}|jm\rangle.$ (65)
Using Eq. (62), we finally obtain
@@
$\theta_{\text{G}}(t)=\arg\left[\frac{2\cos\left(\frac{\Theta(t)}{4}\right)\cosh(\beta\omega_{0})+1}{2\cosh(\beta\omega_{0})+1}\right].$ (66)
Figure 4: The IGP $\theta_{\text{G}}$ as a function of $T$ and $\Theta_{f}$ for the $j=1$ one-axis SSS, where $\Theta_{f}\in[0,4\pi]$. $\theta_{\text{G}}$ exhibits discontinuities with respect to both $\Theta_{f}$ and $T$. At finite temperatures, $\theta_{\text{G}}$ always experiences a $\pi$-jump when crossing $\Theta_{f}^{c}=4\arccos\left(-\operatorname{sech}\left(\frac{\omega_{0}}{T}\right)/2\right)$, indicating a geometric phase transition. When $\Theta_{f}\in(2\pi,8\pi/3)$, there also exists a finite-temperature geometric phase transition as temperature varies.
To visualize the behavior of the IGP, we show in Fig. 4 $\theta_{\text{G}}$ as a function of temperature $T$ and the final squeezing parameter $\Theta_{f}$. Here $\Theta_{f}\equiv\Theta(t_{\text{final}})$ denotes the value of $\Theta$ at the end of the evolution governed by $\tilde{S}(t)$. As the system crosses a critical value $\Theta_{f}^{c}$ at finite temperatures, the IGP undergoes a sudden jump from 0 to $\pi$. In the zero-temperature limit, the critical squeezing parameter approaches $\Theta_{f}^{c}(T\to 0)=2\pi$ while in the high-temperature limit it tends to $\Theta_{f}^{c}(T\to\infty)=\frac{8\pi}{3}$. Similar to the previous example involving the CSS, the IGP according to the density matrix $\rho(t_{\text{final}},T_{c}^{-})$ remains in phase with $\rho(0,T_{c}^{-})$, whereas that of $\rho(t_{\text{final}},T_{c}^{+})$ becomes out of phase with $\rho(0,T_{c}^{+})$ for the $j=1$ one-axis SSS. In general, the relation between $\Theta^{c}_{f}$ and $T_{c}$ for the $j=1$ one-axis SSS can be determined from the solutions of
$2\cos\left(\frac{\Theta(t)}{4}\right)\cosh(\beta\omega_{0})+1=0,$ (67)
as implied by Eq. (66). Explicitly, $\Theta^{c}_{f}=4\arccos\left(-\frac{1}{2}\operatorname{sech}\left(\frac{\omega_{0}}{T_{c}}\right)\right).$ Therefore, the IGP of the one-axis SSS can exhibit finite-temperature geometric phase transitions, analogous to those observed in the CSS. This highlights the rich and intriguing geometric structures inherent in the SSS.
@@ IV.2 Two-Axis Squeezing
Here we analyze the two-axis squeezing scenario and analyze its mixed-state geometric phases exemplified by the Uhlmann phase and IGP.
@@ IV.2.1 Uhlmann Phase
Figure 5: Uhlmann phase $\theta_{U}$ of the $j=1$ two-axis SSS as a function of temperature $T$ on semi-log scale. The smooth curve indicates the absence of a transition.
@@
Figure 6: (Top panel) Contour plot of the IGP $\theta_{\text{G}}$ of the $j=1$ two-axis SSS as a function of $T$ and $\theta_{f}$, where $\theta_{f}\in[0,\frac{3\pi}{4}]$. $\theta_{\text{G}}$ is continuous with respect to both $\theta_{f}$ and $T$ (apart from a jump at $T=0$). (Bottom panels) $\theta_{\text{G}}$ vs $T$ at $\theta_{f}=\pi/2$, where $\theta_{\text{G}}$ exhibits no jumps (left), and $\theta_{\text{G}}$ vs $\theta_{f}$ at $T=0$, where there is a $\pi$-jump at $\theta_{f}\approx 1.33$rad (right).
@@ IV.2.2 The IGP
where all diagonal elements are zero. Therefore, the IGP parallel-transport condition (20) is equivalent to
@@ V Implications
Since the CSS and SSS have been realized in condensed matter or atomic, molecular, and optical systems [54, 55, 71, 72, 73], experimental generation and measurement of the mixed-state geometric phases may become feasible in the future. However, imposing the parallel-transport condition for the IGP or Uhlmann phase requires precise controls of the systems and the ancilla representing the environment, which can be challenging for natural systems in laboratories.
On the other hand, the rapid development of quantum computers opens possibilities of simulating complex quantum systems. By using an ancilla qubit acting as the environment and coupling it to a system qubit, Ref. [74] shows that the Uhlmann phase of a spin-$1/2$ system can be simulated on a quantum computer by using two engineered Hamiltonians to respectively evolve the system and ancilla simultaneously. Since a spin-$j$ object may be decomposed as a collection of $2j$ spin-$1/2$ objects, one may use $2j$ qubits as a system of a spin-$j$ system coupled to an ancilla system of $2j$ qubits to realize the CSS or SSS coupled to the environment. It then follows a series of technical challenges to construct suitable engineered Hamiltonians for the system and ancilla in order to generate the corresponding IGP or Uhlmann phase when the corresponding parallel-transport condition is imposed during the evolution. After the targeted geometric phase is generated on a quantum computer, one may follow Ref. [74] to couple additional measurement qubits to the composite system and perform state tomography to extract the phase accumulated during the constrained evolution.
@@ VI Conclusion
We have derived the Uhlmann phase and IGP for the CSS and SSS. Selected examples are fully solved to demonstrate interesting properties of those mixed-state geometric phases. For the $j=3/2$ CSS, the Uhlmann phase exhibits topological phase transitions characterized by abrupt, quantized jumps while the IGP shows geometric phase transitions with discontinuous jumps as temperature varies. For the $j=1$ one-axis SSS, both Uhlmann phase and IGP display similar temperature-dependent transitions. In contrast, the $j=1$ two-axis SSS shows different behavior as both Uhlmann phase and IGP vary smoothly with temperature, suggesting the absence of temperature-induced phase transitions. These findings highlight the rich interplay between temperature, spin properties, and geometric phases, and offer insights into topological and dynamic behavior of quantum systems at finite temperatures. Furthermore, our results pave the way for potential experimental explorations and numerical simulations to probe interesting spin states with implications for quantum metrology, quantum information processing, and the study of topological and geometric quantum phase transitions.
@@ ACKNOWLEDGEMENTS
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X.Y.H. was supported by the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611). C.C.C. was supported by the NSF (No. PHY-2310656) and DOE (No. DE-SC0025809).
@@ Appendix A Correspondence between Uhlmann phase and Berry phase of CSS as $T\rightarrow 0$
where $J_{z}|j,-j\rangle=-j|j,-j\rangle$ has been applied. This result precisely reduces to the Berry phase of the ground state $|j,-j\rangle$ [57]. More specifically, when $j=\frac{3}{2}$ and $\theta=\frac{\pi}{2}$, $|\zeta|=1$ and thus $\theta_{U}=\mathrm{i}\frac{3}{4}{\oint}_{\gamma}(-2\mathrm{i}\tan^{2}\frac{\theta}{2})\mathrm{d}\phi=3\pi\equiv\pi\mod 2\pi$, which confirms the result shown in Fig. 1 as $T\rightarrow 0$.